OpenStudy (anonymous):

Determine the number and type of solutions for x^2 + 2x + 1= 0. Answer A. 1 Rational Double B. 2 Rational C. 2 Irrational D. 2 Complex Determine the number and type of solutions for x^2 + 2x + 1= 0. Answer A. 1 Rational Double B. 2 Rational C. 2 Irrational D. 2 Complex @Mathematics

OpenStudy (anonymous):

test discriminant = (2)^2 - 4* 1 * 1 = O

hero (hero):

x^2 + x + x + 1 = 0 x(x+ 1)+1(x+1) =0 (x+1)(x+1) = 0 (x+1)^2 = 0 x+1 = 0 x = -1

hero (hero):

A

OpenStudy (anonymous):

a - RATIONAL double

OpenStudy (anonymous):

Thanks :)

OpenStudy (anonymous):

yw

hero (hero):

Como se dice "you're welcome" en espanol?

OpenStudy (anonymous):

Te nada lol

hero (hero):

It's nothing

OpenStudy (anonymous):

Yeah..